The sides of ne triangle have lengths 8, 15, and 17. The perimeter of a similar triangle is 120. Find the lengths of the sides of the second triangle.
Answer by checkley71(8403)

8+15+17=40
120/40=3 this is the multiplier for all sides. THUS:
8*3=24
15*3=45
17*3=51
proof:
24+45+51=120
120=120